OpenStudy (anonymous):

Auditions for a new, upcoming band in Chicago were held with 90 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 30, how many women auditioned? 50 40 60 30

OpenStudy (anonymous):

30

OpenStudy (anonymous):

67

OpenStudy (anonymous):

I mean 60 sorry :P

OpenStudy (anonymous):

thanks pac i chose 60

OpenStudy (anonymous):

tyto leave me alone

OpenStudy (lgbasallote):

i can randomly say 40...doesnt make me any more wrong than any of these people here ;)

OpenStudy (anonymous):

^LOL

OpenStudy (anonymous):

^lolololol

OpenStudy (anonymous):

your so right.....PAC SHOW ME HOW UGOT THAT ANSWER

OpenStudy (anonymous):

I did it by magic. ;)

OpenStudy (anonymous):

I am boss :D

OpenStudy (anonymous):

50. Thats my final answer!

OpenStudy (anonymous):

tyto get off my question

OpenStudy (anonymous):

right now

OpenStudy (anonymous):

Make me. >.>

OpenStudy (anonymous):

thats impossible <.<

OpenStudy (anonymous):

Exactly! :D

OpenStudy (lgbasallote):

let x = male y = female x + y = 90 -> x = 90-y 2x - 3y = 30 2(90-y) - 3y = 30 180 - 2y - 3y = 30 180 - 5y = 30 180 - 30 = 5y 150 = 5y y= 30

OpenStudy (lgbasallote):

see? i told you they were wrong :S

OpenStudy (anonymous):

I was so right. And you weren't even thankful Jack.

OpenStudy (anonymous):

I WAS RIGHT :D

OpenStudy (lgbasallote):

@Pac1f1cIslander 60 is the number of men

OpenStudy (lgbasallote):

30 is the number of female

OpenStudy (anonymous):

And Men are awesome! :D

OpenStudy (anonymous):

Oh wait..I was actually right?

OpenStudy (anonymous):

*facepalm*

OpenStudy (lgbasallote):

this is why it is customary here to show solution haha

OpenStudy (anonymous):

lgba i noticed your not an ambassador anymore